a room is 11ft by 19ft it needs to be painted the ceiling is 8ft above the floor there are 2 windows in the room each window is 3ft by 4ft the door is 2.5ft by 6.5ft. Find the area of the walls and ceiling that need to be painted?
Answer by richwmiller(17219)

ceiling 11*19=c
2 walls 2*8*11=w
2 walls 2*8*19=s
2 windows 2*3*4=n
1 door 2,5*6.5=d
c+w+s-n-d=area